How Do You Graph y=600x ?
Found 2 solutions by blwinbbbles, mathhater404:
Answer by blwinbbbles(106)

I am assuming you know how to graph using x and y cordinates...
so if y=600x then you would just pick any "x" so if x = 0 then
y=600(0) so y=0 ...now you know that when x = 0 then y = 0 so that cordinate would be (0,0)
if x = 1 then y = 600(1) so y = 600
so when x = 1, then y = 600 so the second cordinate is (1,600)
now you can graph the two points and draw a linel through the two points. This would represent a graph of y= 600(x)

Answer by mathhater404(3)

Your graphing coordinates are off y should equal a number so the tutors awnser is wrong because he could not have a y if there never was a y but the x numbers are right